OpenStudy (anonymous):

Solve: x2 – 16x + 15 = 0 a) x = -1, x = -15 b) x = 1, x = 15 c) x = 3, x = 5 d) x = -3, x = -5

OpenStudy (anonymous):

x^2-16x +15=0 (x-1)(x-15)=0 x=1 or x =15 therefore, b is the ans